Find the consumers surplus at Q = 2 for the demand function P=41-902 cs (Simplify your answer.)
Answer #1

- The demand function p=41-90² . at Q=2, . P2 HI- 9 (2)2- 41-(9x4) Demand function at Q=0 =) at p=0 a) = 5 P= 41-90² P = 41 .Huce consues surples - = - Alea Of A ABC Ž x base x height { X (2-0) x (41-5) = 36 Hrnce es= (s = 1 Hince 36 CS Scanned with